OpenStudy (anonymous):

Can someone take a look at this? "Divide 1/-5+2i answer as a number in complex standard form" bear in mind i=sqrt-1

OpenStudy (anonymous):

\[\frac{1}{-5 + 2i}\]

OpenStudy (anonymous):

\[\frac{1}{-5 +2i} \times \frac{5 +2i}{2 +2i}\]

OpenStudy (anonymous):

its 5in denominator not 2 => 5 + 2i

OpenStudy (anonymous):

So ..when we rationalize we get -(5 + 2i) -5 2i --------= ---- - ----- 25 +4 29 29

OpenStudy (anonymous):

Thanks for helpin with this.. this one has been gettin to me